# Rethinking Neural Combinatorial Optimization for Vehicle Routing Problems with Different Constraint Tightness Degrees

### v2

@@ 7.1 NCO on VRPs with Fixed Constraint Tightness
NCO methods primarily learn construction heuristics to arrange input nodes into feasible solutions. The pioneering Ptr-Net [18] introduces a Recurrent Neural Network (RNN)-based architecture to solve TSP through supervised learning. In contrast, Bello et al. [25] employs reinforcement learning to eliminate dependency on labeled solutions. Subsequent work by Nazari et al. [26] extends this framework to solve CVRP. A significant advancement comes with the Attention Model [13], which adopts the Transformer architecture [27] to achieve state-of-the-art performance across various VRPs with up to 100 nodes. This inspires numerous AM-based enhancements [6, 28, 22, 29, 30, 31, 32, 33] that narrow the performance gap with classical heuristic solvers. However, these methods share a critical limitation: they are typically trained on VRP instances with fixed constraint tightness. For example, models trained on CVRP100 instances typically adopt a fixed capacity of C=50 [6, 34, 14, 15, 20, 21, 23, 17]. This training paradigm may overfit the NCO models on training instances with certain specific constraint tightness, restricting their adaptability to other cases. As a result, these models suffer severe performance degradation when applied to instances with out-of-domain constraint tightness. Our work systematically investigates the underlying causes of this limitation and provides new insights to tackle it.
@@ Conclusion
Appendix A Detailed Formulation of Attention Layer
A.1 Attention Layer
$X^{(\ell)}=\mbox{Norm}(\mbox{FF}(\hat{X})+\hat{X}).$
For simplicity, the entire process can be represented by
$X^{(\ell)}$ $=\mbox{AttnLayer}(X^{(\ell-1)},Y^{(\ell-1)}).$ (6)
A.2 Multi-head Attention
We first describe the single-head attention function as follows:
$\mbox{head}_{i}$ $=\mbox{Attn}_{i}(X,Y),$
A.3 Feed-Forward Layer
$\mbox{FF}(X)=\mbox{max}(0,XW_{1}+b_{1})W_{2}+b_{2},$ (9)
Appendix B Quantification of Problem Similarity
Problem Similarity between CVRP and OVRP/TSP
Table 6: Problem similarity analysis across capacity constraints. Transferring rules: OVRP→CVRP adds return trips to the depot; CVRP→TSP removes capacity constraints, and TSP→CVRP reintroduces them. We detail solution transformation in Appendix C.
CVRP100
C10 C50 C400 C500
OVRP cost 30.88 9.84 7.49 7.49
CVRP cost 59.77 15.55 8.04 7.87
TSP cost 7.80 7.80 7.80 7.80
OVRP→CVRP cost 60.67 17.25 9.78 9.74
CVRP→OVRP cost 31.18 11.33 7.76 7.70
TSP→CVRP cost 70.47 17.91 8.78 7.98
CVRP→TSP cost 13.17 10.68 7.96 7.83
Similarity CVRP-OVRP 97.5% 75.6% 75.5% 74.1%
Similarity CVRP-TSP 25.5% 53.5% 88.9% 98.2%
Problem Similarity between CVRP and CVRPTW
CVRP100
$\alpha$ = 0.2 1.0 3.0 5.0
CVRP cost 15.51 15.51 15.51 15.51
CVRPTW cost 33.77 24.42 15.81 15.55
CVRP→CVRPTW cost 59.69 38.89 18.89 16.08
CVRPTW→CVRP cost 24.42 21.24 15.80 15.55
Similarity CVRP-CVRPTW 9.9% 25.7% 79.0% 96.3%
Appendix C Details for Solution Transformation
C.1 CVRP-TSP
  • Converting CVRP Solution to TSP Solution
    1. 1.
    2. 2.
    1. 1.
      Insert a depot node after the current node.
    2. 2.
      Reset the cumulative demand value to zero.
    3. 3.
      Start visiting the next node from the depot again.
C.2 CVRP-OVRP
C.3 CVRP-CVRPTW
  • Converting CVRP Solution to CVRPTW Solution
    1. 1.
    2. 2.
    3. 3.
    Third, the path direction optimization is performed. We perform bidirectional time window-based path splitting for each sub-tour: Calculate the total travel distance of split sub-tours performing the time window-based path splitting on the sub-tour for both the forward and reverse directions, separately. And finally choose the sub-tour with the shorter distance.
  • Converting CVRPTW solution to CVRP Solution
    1. 1.
    2. 2.
Appendix D Batch-level vs. Instance-level Strategy for Assigning Constraint Tightness
Table 8: Batch-level vs. Instance-level random $C$ for model training.
CVRP100
C10 C50 C100 C200 C300 C400 C500
Gap Gap Gap Gap Gap Gap Gap
Batch-level random $C$ 1.99% 4.66% 5.02% 2.46% 1.06% 0.89% 0.94%
Instance-level random $C$ 1.54% 3.82% 3.67% 1.49% 0.87% 0.75% 0.87%
Appendix E Quantification of Time Window Constraint Tightness for CVRPTW
CVRPTW
Following [42, 43, 44], we define the CVRPTW as follows. Building upon CVRP, CVRPTW introduces a time window $[e_{i},l_{i}]$ for each customer node $i$, representing the earliest and latest times at which customer $i$ can be visited. Additionally, each node has a service time $s_{i}$, which represents the duration a vehicle must remain at the node upon arrival. The depot has a time window $[e_{0},l_{0}]$ but no service time. Each vehicle departs from the depot at time $e_{0}$ and must return before $l_{0}$. Let $i$ and $j$ be two consecutive nodes on sub-tour $r$, and let $b_{i}^{r}$ denote the time at which the vehicle on sub-tour $r$ arrives at node $i$. The departure time from node $i$ is given by $b_{i}^{r}+s_{i}$, and the arrival time at the next node $j$ is calculated as: $b_{j}^{r}=\operatorname{max}(e_{j},b_{i}^{r}+s_{i}+t_{ij})$, where $t_{ij}$ is the travel time from node $i$ to node $j$, equal to the Euclidean distance between the two points. The calculation of arrival time indicates that if the vehicle arrives at node $j$ earlier than its start time $e_{j}$, it must wait until $e_{j}$ to begin service. If the vehicle arrives at node $j$ later than $l_{j}$, the sub-tour becomes infeasible. The objective of CVRPTW is to minimize the total Euclidean distance of all sub-tours.
Time Window Tightness
$\delta=\frac{l_{i}-e_{i}}{2}\cdot(1-\alpha),$ (11)
$e_{i}^{\prime}=\max(e_{i}+\delta,0)$
$l_{i}^{\prime}=l_{i}-\delta,$
Appendix F Implementation Details of Multi-expert Module in POMO
POMO model
$h_{c}^{\prime}$ $=\mbox{MHA}(h_{c},H_{t}),$ (12)
where $C_{l}$ is a constant.
POMO model with Multi-expert Module
We substitute the decoder in POMO with the multi-expert module. The multi-expert module is generally the same as the one presented in Section 4.3, but different in the structure of the expert layer. Since the POMO model is typically trained using reinforcement learning, the expert layer cannot be too heavy to avoid excessive memory usage and long training time. Therefore, built on the original decoder, each expert layer only adds one feed forward layer into it. We formulate the expert layer as follows:
$h_{c}^{\prime}$ $=\mbox{MHA}(h_{c},H_{t})$ (13)
$\hat{h}_{c}$ $=\mbox{FF}(h_{c}^{\prime})+h_{c}^{\prime}$
$P$ $=\mbox{Softmax}\mbox{(}C_{l}\cdot\mbox{tanh}\frac{(\hat{h}_{c})^{\intercal}H_{t}}{\sqrt{d_{h}}})$
Appendix G Broader Impacts
This research contributes to the field of neural combinatorial optimization by revisiting the training setting of existing NCO models, finding that models trained under these settings may overfit certain specific constraint tightness, and proposing a simple yet efficient training scheme and a multi-expert module to tackle the overfitting issue. We believe our insights and proposed methods can provide valuable insights and inspire follow-up works to explore more robust neural methods designed for solving VRPs. As general learning-based methods for solving VRPs, the proposed methods do not have any specific potential negative societal impact.
Appendix H Licenses
The licenses for the codes used in this work are listed in Table 9.
Table 9: Licenses for codes used in this work
Resource Type Link License
HGS [16] Code https://github.com/chkwon/PyHygese MIT License
POMO [6] Code https://github.com/yd-kwon/POMO MIT License
BQ [14] Code https://github.com/naver/bq-nco CC BY-NC-SA 4.0
ELG [17] Code https://github.com/gaocrr/ELG MIT License
POMO-MTL [23] Code https://github.com/FeiLiu36/MTNCO MIT License
MDAM [22] Code https://github.com/liangxinedu/MDAM MIT License
@@ 18. Answer: [Yes]
  19. Justification: We have provided detailed implementation details and experiment settings in Section F and Section 5, respectively, to ensure reproducibility.
@@ 20. Guidelines:
  23. Answer: [No]
  24. Justification: We will release the coder once upon acceptence.
@@ 45. Guidelines:
  48. Answer: [Yes]
  49. Justification: The proposed learning-based method for solving VRPs is not anticipated to have any specific negative societal impacts. We provide a detailed discussion in Appendix G.
@@ 58. Answer: [Yes]
  59. Justification: We have included the licenses of baseline solvers and datasets in Section H.

### v3

@@ 7.1 NCO on VRPs with Fixed Constraint Tightness
NCO methods primarily learn construction heuristics to arrange input nodes into feasible solutions. The pioneering Ptr-Net [27] introduces a Recurrent Neural Network (RNN)-based architecture to solve TSP through supervised learning. In contrast, Bello et al. [30] employs reinforcement learning to eliminate dependency on labeled solutions. Subsequent work by Nazari et al. [31] extends this framework to solve CVRP. A significant advancement comes with the Attention Model [18], which adopts the Transformer architecture [32] to achieve state-of-the-art performance across various VRPs with up to 100 nodes. This inspires numerous AM-based enhancements [9, 33, 22, 34, 35, 36, 37] that narrow the performance gap with classical heuristic solvers. However, these methods share a critical limitation: they are typically trained on VRP instances with fixed constraint tightness. For example, models trained on CVRP100 instances typically adopt a fixed capacity of C=50 [9, 38, 19, 20, 24, 26, 25, 23]. This training paradigm may overfit the NCO models on training instances with certain specific constraint tightness, restricting their adaptability to other cases. As a result, these models suffer severe performance degradation when applied to instances with out-of-domain constraint tightness. Our work systematically investigates the underlying causes of this limitation and provides new insights to tackle it.
@@ Conclusion
Acknowledgments and Disclosure of Funding
This work was supported in part by the National Natural Science Foundation of China (Grant 62476118), the Natural Science Foundation of Guangdong Province (Grant 2024A1515011759), the Guangdong Science and Technology Program (Grant 2024B1212010002), and the Center for Computational Science and Engineering at Southern University of Science and Technology.
@@ 18. Answer: [Yes]
  19. Justification: We have provided detailed implementation details and experiment settings in Section G and Section 5, respectively, to ensure reproducibility.
@@ 20. Guidelines:
  23. Answer: [Yes]
  24. Justification: We have released the code, the link to which is shown in the abstract.
@@ 45. Guidelines:
  48. Answer: [N/A]
  49. Justification: The proposed learning-based method for solving VRPs is not anticipated to have any specific negative societal impacts.
@@ 58. Answer: [Yes]
  59. Justification: We have included the licenses of baseline solvers and datasets in Section P.